OpenStudy (anonymous):

what's the trig identity for sin^2(x) = ? so that I can integrate it.

OpenStudy (psymon):

(1-cos2x)/2

OpenStudy (anonymous):

yes. thank you!

OpenStudy (psymon):

:3

OpenStudy (anonymous):

recall $$\cos2x=\cos^2 x-\sin^2x=1-\sin^2x-\sin^2x=1-2\sin^2 x$$Solving for \(\sin^2 x\) we have:$$2\sin^2 x=1-\cos 2x\\\sin^2 x=\frac12(1-\cos 2x)$$

OpenStudy (anonymous):

true. Unfortunately this is part of a much bigger problem and I didn't want to go to all that trouble :)